Explain the steps you would take to solve the problem.
7 + x/5 = -4

Answer by Theo(13342)

the equation is 7 + x/5 = -4
subtract 7 from both sides of the equation to get:
7 + x/5 - 7 = -4 - 7
combine like terms to get:
x/5 = -11
multiply both sides of the equation by 5 to get:
x/5 * 5 = -11 * 5
simplify to get:
x = -55


to confirm, replace x in the original equation with -55 to get:
7 + x/5 = -4 becomes 7 + -55/5 = -4
simplify to get:
7 - 11 = -4
combine like terms to get:
-4 = -4
this confirms the solution is correct.